# I/O in Machine Learning Applications on HPC Systems: A 360-degree Survey

### v1

@@ 5.2. Proposed I/O Optimization Techniques
With the increasing complexity of ML models, several methods have been proposed to enable the training of models that do not fit into accelerator or even clustered accelerator memory. Previous methods aiming to reduce memory usage during training include materialization and offloading. Materialization deletes some activations during the forward pass, which are then recomputed during the backward pass trading computation for memory. Offloading moves activations from the limited GPU memory to CPU memory during the forward pass trading data movement for memory (Beaumont et al., 2022). The limited HPC bandwidth, in addition to the increased communication caused by the offloading of the model to storage, is a large challenge for these strategies. Rajbhandari et al. (Rajbhandari et al., 2021), propose ZeRO-Infinity to train large models by leveraging GPU, CPU, and NVMe memory. ZeRO-Infinity is fully integrated with PyTorch. Jang et al. (Jang et al., 2024) propose Smart-Infinity to train large models with storage offloading. Smart-Infinity is also integrated with PyTorch and takes advantage of storage devices to form an extended memory hierarchy. In contrast to ZeRO-Infinity, they use near-storage accelerators to reduce I/O pressure and increase training speeds. In addition, gradient compression/decompression is used to reduce the storage congestion incurred from model offloading. Sun et al. (Sun et al., 2022) propose STRONGHOLD for large model training. They offload model training from the GPU to the CPU; however, the minimum amount of data to be kept on the GPU to minimize GPU memory usage is dynamically determined at runtime. In addition, they use prefetching to load layers of the model in parallel with GPU computation.
@@
In GPU architectures, storage is typically organized into tiers, comprising high-bandwidth memory (HBM) and on-chip SRAM, facilitating efficient data access and processing (Zhu et al., 2018). Doa et al. (Dao et al., 2022) propose FlashAttention which uses an I/O aware exact attention algorithm to reduce the number of memory read and writes that occur between HBM and SRAM on GPU devices. It efficiently moves data between the GPU storage tiers resulting in faster training speeds while maintaining the same level of accuracy as traditional attention algorithms. FlashAttention has been adopted into several machine learning frameworks such as PyTorch’s nn.Transformer, Microsoft’s DeepSpeed (Rasley et al., 2020), and NVIDIA’s MEgatron-LM (Shoeybi et al., 2020). In addition it yields the fastest in MLPerf (Mattson et al., 2020) BERT training on cloud instances in MLPerf training (June 2022) and MLPerf training (November 2022).
@@
FFCV (Leclerc et al., 2023) is a library aiming to speed up the data loading and processing pipeline while easily integrating into existing PyTorch applications. FFCV offers a custom file format that can store arbitrary data modalities and enable non-sequential data reading. It stores all samples in a single file but splits the dataset into pages, which trades space for more efficient data reads. If the dataset can fit into memory, it will be cached, and every read of a sample after the first will be read from memory instead of the underlying storage system. In contrast, if the dataset does not fit into memory, FFCV enables efficient data loading by using the PRNG to prefetch data samples. In addition, it uses quasi-random data shuffling in which a permutation of the data samples is loaded into a memory buffer large enough for pages of the dataset. The model then selects from batches existing within the buffer. This reduces the number of accesses to the underlying storage system. FFCV aims to eliminate unnecessary memory allocations by having operations in the data loading and processing pipeline declare memory requirements so that memory allocation occurs only once before each epoch. It also uses just-in-time (JIT) compilation, where possible, during the data processing pipeline to convert Python code to machine code. Since the machine code is no longer under the constraints of the Python interpreter, it enables multi-threading, which by extension enables collaborative reading and writing to and from memory instead of expensive primitives, meaning threads can share batches of data. Threads also enable data preparation, such as data copying and augmentations, to be run in parallel because they share the same CUDA context.
@@
Deep Lake (Hambardzumyan et al., 2022) is a data lake for deep learning. It can be used with object storage systems such as AWS S3 (Amazon Web Services, nd), Google Cloud Storage (GCS) (Google Cloud, nd), POSIX file systems, or local in-memory storage. Deep Lake has a Tensor Query Language that extends the SQL Parser from Hyrise (Dreseler et al., 2019). This enables multidimensional array operations and deeper integration with the data lake, including indexing, slicing of arrays, and a large set of convenience functions. There is also a data visualizer made available through a web interface, allowing for the visualization of large-scale data. The Deep Lake streaming data loader delegates data fetching to C++ processes to avoid Python’s GIL. Deep Lake also has a version control system that keeps track of different branches of the dataset in the same storage. This enables machine learning engineers to version their datasets for further querying. Another distinction is Deep Lake’s ability to store unstructured data with its metadata in learning-native columnar format, enabling fast data streaming.
@@
With the growing size of AI models, efficient checkpointing methods are needed to ensure scalability. Nicolae et al. (Nicolae et al., 2020) propose DeepFreeze, an asynchronous checkpointing technique built using Very Low Overhead Checkpoint-Restart (Veloc) (Nicolae et al., 2019). Veloc is a lightweight concurrency-optimized checkpointing library targeted at delivering high-performance checkpointing on top of heterogeneous storage hierarchies. DeepFreeze is a transparent checkpointing solution for synchronous data parallel training with layer-wise model parallelism. To increase checkpointing speeds, an asynchronous multi-level checkpointing approach is used in which a local copy of the learning model is captured and persisted in lightweight (i.e. node-local storage of neighboring nodes) and heavy (i.e. PFS) storage to block training for the least amount of time possible. To reduce the overhead of model serialization, a compact binary format is used that leaves out details such as tensor labels. The model’s layers are sliced into shards equal to the number of MPI ranks. The ranks then write checkpoints for their partition of the model to local storage. This allows for models to be checkpointed in parallel, avoiding the stragglers problem. Anthony et al. (Anthony and Dai, 2021) also propose local process model snapshots, which are then stored in node-local storage. Their method uses the SCR-Exa library (Dai, 2022), which is a scalable checkpointing tool for HPC applications. There are two restart methods (cold and hot) available. Cold restart attempts to restart the application using the checkpointing caching within the same allocation of nodes. The hot restart tries to replace faulty nodes with idle spare nodes. Check-N-Run (Eisenman et al., 2022) is a checkpointing system for deep-learning recommendation models developed by Facebook. It uses two techniques to address size and bandwidth: differential checkpointing, where only a fraction of the model changes per iteration, and quantization techniques to reduce checkpoint size. The checkpoint system is implemented under the PyTorch framework. Check-N-Run keeps training accuracy as a priority as they state an accuracy loss as small as is unacceptable.
@@ 6. Future Research Directions of I/O for ML
While the ML and AI surge is rushing forward driven by rapidly increasing data collection and application use cases, data management and I/O software have been significantly lagging, especially in the HPC domain. In this section, we identify numerous data management and I/O areas that must be improved for ensuring continued progress and innovation in the training and deployment of ML models.
@@
When preparing data from remote storage locations such as data warehouses, it may be necessary to store intermediate forms of the dataset after preprocessing. Further research could involve an I/O middleware that transparently stores intermediate forms of the dataset in clustered memory or the underlying storage system, depending on the dataset size. By extending the memory to the underlying storage, additional questions arise, such as how data movement between memory and storage can be optimized for high throughput and low latency and what caching strategies should be used to ensure efficient data access.
@@
With the exponential growth of data in machine learning, ensuring data quality is essential for the accuracy and reliability of models (Jain et al., 2020). One step towards ensuring high-quality data is to use data warehouses because they often provide data governance tools. However, scientists who use HPC systems for high I/O throughput could take a significant performance hit if limited by the bandwidth of the external network. This motivates the need for tools that can aid in obtaining and monitoring high-quality datasets. These tools could offer several data management services, such as integrity constraints, versioning, provenance, and quality assessments while being scalable and supporting a variety of data modalities. A recent survey on metrics for data readiness for AI has exposed lack of metrics focused on quality, bias, and FAIR (Findability, Availability, Interoperability, and Reusability) principle compliance (Hiniduma et al., 2024).
@@
Due to the rise of the cloud-based storage being used by ML workloads, methods of efficiently streaming data from multiple data sources are needed. In addition, multi-modal models may need to stream data from several data sources, each with varying data modalities. Future research could involve an I/O middleware capable of concurrently optimizing I/O from multiple data sources. The middleware would take into account the varying latencies between the data sources and use methods such as sample caching and sample prefetching to ensure optimal performance. This would involve a cache eviction strategy in which eviction is based on the dynamic latency of the data sources. In addition, prefetch scheduling could also occur dynamically, ensuring that samples from high-latency data sources have a higher priority than low-latency data sources. It would also take into account the sample size when performing these optimizations to ensure the optimal balance between latency and data size.
@@
GPUs have become increasingly popular due to the rise of ML. This is due to their high level of parallelism when doing mathematical computations and, in particular, matrix operations (Steinkraus et al., 2005). This is possible because GPUs, unlike CPUs, contain thousands of cores. This has led to the development of ML domain-specific processing units, such as tensor processing units (TPU)s by Google (Jouppi et al., 2017). These are specialized processors specifically built for ML workloads. With the wide range of hardware available, future research is needed to determine how data movement and storage techniques should be used effectively across the diverse range of hardware available. Questions for future research include where to place and process samples to prioritize processing units and exploit their contrasting abilities effectively, how to efficiently move data between the diverse range of processing units, and how to provide transparent and reasonable solutions for developers facing these challenges.
Removing redundant I/O and processing during inference
A strategy proposed by Xie et al. (xie et al., 2022) discussed in Section 5 aims to reduce redundant I/O and computation in the training phase by caching intersecting samples after preprocessing, which are used by multiple processes. However, during concurrent model deployment, dataset reading is commonly not done in a uniformly random fashion, causing their caching strategy to be non-applicable. In addition, their work does not consider storing intersecting samples at varying levels of the storage hierarchy depending on how many processes intersect at a given sample. Future work could involve devising a caching solution that works for less restrictive access patterns and considers the degree of overlap among processes at a given sample to determine what storage level to use for caching, increasing the efficiency and scalability of concurrent model execution.
@@
As shown in Fig. 1, HPC I/O and ML I/O stacks have been organically evolving for faster implementation of the ML algorithms although there is significant scope for higher performance. As the ML algorithms settle, improving efficiency of data access in running various applications both on HPC systems as well as cloud computing environments need to take priority. ML application development should also target obtaining portability of efficiency on different platforms. Achieving this will require efforts in benchmarking, evaluation of performance with various programming tools, and developing platform-specific optimization strategies.
@@ 7. Conclusion
Data plays a pivotal role in machine learning, impacting the learning process and significantly affecting model performance and accuracy. With a wide array of file formats available, each offering distinct I/O optimizations, machine learning applications must carefully select the format that best suits their needs. Additionally, the modality of the dataset itself greatly influences the choice of file formats, affecting how datasets are accessed and stored throughout the machine learning lifecycle.
@@
Modern ML frameworks commonly offer several options to optimize data loading and processing, such as parallel data preparation, dataset streaming, and sample prefetching, to optimize the processing and loading of batches of samples. However, caching solutions that work with datasets that do not fit into node local memory are not widely available in ML frameworks such as PyTorch and TensorFlow, often leading to sample accesses always reaching the underlying storage system which may lead to an I/O bottleneck. Recent advances toward minimizing I/O time include specialized shuffling techniques, I/O schedulers, and domain-specific file formats.
@@ Acknowledgements.
- Abadi et al. (2008) Daniel J. Abadi, Samuel R. Madden, and Nabil Hachem. 2008. Column-stores vs. row-stores: how different are they really?. In Proceedings of the 2008 ACM SIGMOD International Conference on Management of Data (Vancouver, Canada) (SIGMOD ’08). Association for Computing Machinery, New York, NY, USA, 967–980. https://doi.org/10.1145/1376616.1376712
@@
- Adolf et al. (2016) Robert Adolf, Saketh Rama, Brandon Reagen, Gu-Yeon Wei, and David M. Brooks. 2016. Fathom: reference workloads for modern deep learning methods. 2016 IEEE International Symposium on Workload Characterization (IISWC) (2016), 1–10. https://api.semanticscholar.org/CorpusID:1809844
@@
- Ali et al. (2014) Peshawa Jamal Muhammad Ali, Rezhna Hassan Faraj, Erbil Koya, Peshawa J Muhammad Ali, and Rezhna H Faraj. 2014. Data normalization and standardization: a technical report. Mach Learn Tech Rep 1, 1 (2014), 1–6.
- Alonso-Monsalve et al. (2021) Saúl Alonso-Monsalve, Alejandro Calderón, Félix García-Carballeira, and José Rivadeneira. 2021. Analyzing the distributed training of deep-learning models via data locality. In 2021 29th Euromicro International Conference on Parallel, Distributed and Network-Based Processing (PDP). 117–121. https://doi.org/10.1109/PDP52278.2021.00026
@@
- Apache Software Foundation (2024) Apache Software Foundation. 2024. Apache Spark. https://spark.apache.org/.
@@
- Bez et al. (2021) Jean Luca Bez, Houjun Tang, Bing Xie, David Williams-Young, Rob Latham, Rob Ross, Sarp Oral, and Suren Byna. 2021. I/O Bottleneck Detection and Tuning: Connecting the Dots using Interactive Log Analysis. In 2021 IEEE/ACM Sixth International Parallel Data Systems Workshop (PDSW). 15–22. https://doi.org/10.1109/PDSW54622.2021.00008
@@
- Brown et al. (2020) Tom B. Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel M. Ziegler, Jeffrey Wu, Clemens Winter, Christopher Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever, and Dario Amodei. 2020. Language Models are Few-Shot Learners. arXiv:2005.14165 [cs.CL]
@@
- Chen et al. (2018) Chia-Yu Chen, Jungwook Choi, Daniel Brand, Ankur Agrawal, Wei Zhang, and Kailash Gopalakrishnan. 2018. AdaComp : Adaptive Residual Gradient Compression for Data-Parallel Distributed Training. Proceedings of the AAAI Conference on Artificial Intelligence 32, 1 (Apr. 2018). https://doi.org/10.1609/aaai.v32i1.11728
- Chen et al. (2012) Tianshi Chen, Yunji Chen, Marc Duranton, Qi Guo, Atif Hashmi, Mikko Lipasti, Andrew Nere, Shi Qiu, Michele Sebag, and Olivier Temam. 2012. BenchNN: On the broad potential application scope of hardware neural network accelerators. Proceedings - 2012 IEEE International Symposium on Workload Characterization, IISWC 2012, 36–45. https://doi.org/10.1109/IISWC.2012.6402898
@@
- Chen et al. (2023) Weijian Chen, Shuibing He, Yaowen Xu, Xuechen Zhang, Siling Yang, Shuang Hu, Xian-He Sun, and Gang Chen. 2023. iCache: An Importance-Sampling-Informed Cache for Accelerating I/O-Bound DNN Model Training. In 2023 IEEE International Symposium on High-Performance Computer Architecture (HPCA). 220–232. https://doi.org/10.1109/HPCA56546.2023.10070964
@@
- Chien et al. (2020) Steven W. D. Chien, Artur Podobas, Ivy B. Peng, and Stefano Markidis. 2020. tf-Darshan: Understanding Fine-grained I/O Performance in Machine Learning Workloads. In 2020 IEEE International Conference on Cluster Computing (CLUSTER). IEEE Computer Society, Los Alamitos, CA, USA, 359–370. https://doi.org/10.1109/CLUSTER49012.2020.00046
- Choi et al. (2023a) Hyeonseong Choi, Byung Hyun Lee, Se Young Chun, and Jaehwan Lee. 2023a. Towards accelerating model parallelism in distributed deep learning systems. PLoS One 18, 11 (Nov. 2023), e0293338.
- Choi et al. (2023b) Jong Youl Choi, Massimiliano Lupo Pasini, Pei Zhang, Kshitij Mehta, Frank Liu, Jonghyun Bae, and Khaled Ibrahim. 2023b. DDStore: Distributed Data Store for Scalable Training of Graph Neural Networks on Large Atomistic Modeling Datasets. ACM International Conference Proceeding Series, 941–950. https://doi.org/10.1145/3624062.3624171
- Chowdhury et al. (2019) Fahim Chowdhury, Yue Zhu, Todd Heer, Saul Paredes, Adam Moody, Robin Goldstone, Kathryn Mohror, and Weikuan Yu. 2019. I/O Characterization and Performance Evaluation of BeeGFS for Deep Learning. In Proceedings of the 48th International Conference on Parallel Processing (Kyoto, Japan) (ICPP 2019). Association for Computing Machinery, New York, NY, USA, Article 80, 10 pages. https://doi.org/10.1145/3337821.3337902
@@
- Dau et al. (2018) Hoang Anh Dau, Eamonn Keogh, Kaveh Kamgar, Chin-Chia Michael Yeh, Yan Zhu, Shaghayegh Gharghabi, Chotirat Ann Ratanamahatana, Yanping, Bing Hu, Nurjahan Begum, Anthony Bagnall, Abdullah Mueen, Gustavo Batista, and Hexagon-ML. 2018. The UCR Time Series Classification Archive. https://www.cs.ucr.edu/~eamonn/time_series_data_2018/.
@@
- Devarajan et al. (2021) Hariharan Devarajan, Huihuo Zheng, Anthony Kougkas, Xian-He Sun, and Venkatram Vishwanath. 2021. DLIO: A Data-Centric Benchmark for Scientific Deep Learning Applications. In 2021 IEEE/ACM 21st International Symposium on Cluster, Cloud and Internet Computing (CCGrid). 81–91. https://doi.org/10.1109/CCGrid51090.2021.00018
@@
- Dey et al. (2020) Tonmoy Dey, Kento Sato, Bogdan Nicolae, Jian Guo, Jens Domke, Weikuan Yu, Franck Cappello, and Kathryn Mohror. 2020. Optimizing Asynchronous Multi-Level Checkpoint/Restart Configurations with Machine Learning. In 2020 IEEE International Parallel and Distributed Processing Symposium Workshops (IPDPSW). 1036–1043. https://doi.org/10.1109/IPDPSW50202.2020.00174
@@
- Dryden et al. (2021) Nikoli Dryden, Roman Böhringer, Tal Ben-Nun, and Torsten Hoefler. 2021. Clairvoyant Prefetching for Distributed Machine Learning I/O. In Proceedings of the International Conference for High Performance Computing, Networking, Storage and Analysis (St. Louis, Missouri) (SC ’21). Association for Computing Machinery, New York, NY, USA, Article 92, 15 pages. https://doi.org/10.1145/3458817.3476181

### v2

@@ 6.2. Proposed I/O Optimization Techniques
With the increasing complexity of ML models, several methods have been proposed to enable the training of models that do not fit into accelerator or even clustered accelerator memory. Previous methods aiming to reduce memory usage during training include materialization and offloading. Materialization removes some activations during the forward pass. The removed activations are then recomputed during the backward pass. This strategy trades a larger computational cost for less memory. Offloading moves activations from the limited GPU memory to CPU memory during the forward pass trading data movement for memory (Beaumont et al., 2022). The limited HPC bandwidth, in addition to the increased communication caused by the offloading of the model to storage, is a large challenge for these strategies. Rajbhandari et al. (Rajbhandari et al., 2021), propose ZeRO-Infinity to train large models by leveraging GPU, CPU, and NVMe memory. ZeRO-Infinity is fully integrated with PyTorch. Jang et al. (Jang et al., 2024) propose Smart-Infinity to train large models with storage offloading. Smart-Infinity is also integrated with PyTorch and takes advantage of storage devices to form an extended memory hierarchy. In contrast to ZeRO-Infinity, they use near-storage accelerators to reduce I/O pressure and increase training speeds. In addition, gradient compression/decompression is used to reduce the storage congestion incurred from model offloading. Sun et al. (et al, 2022e) propose STRONGHOLD for large model training. They offload model training from the GPU to the CPU; however, the minimum amount of data to be kept on the GPU to minimize GPU memory usage is dynamically determined at runtime. In addition, they use prefetching to load layers of the model in parallel with GPU computation.
@@
In GPU architectures, storage is typically organized into tiers, comprising high-bandwidth memory (HBM) and on-chip SRAM, facilitating efficient data access and processing (et al, 2018c). Doa et al. (Dao et al., 2022) propose FlashAttention which uses an I/O aware exact attention algorithm to reduce the number of memory read and writes that occur between HBM and SRAM on GPU devices. It efficiently moves data between the GPU storage tiers, resulting in faster training speeds while maintaining the same level of accuracy as traditional attention algorithms. FlashAttention has been adopted into several machine learning frameworks such as PyTorch’s nn.Transformer, Microsoft’s DeepSpeed (et al, 2020g), and NVIDIA’s MEgatron-LM (Shoeybi et al., 2020). In addition, it yields the fastest in MLPerf (et al, 2020d) BERT training on cloud instances in MLPerf training (June 2022) and MLPerf training (November 2022).
@@
FFCV (et al, 2023c) is a library aiming to speed up the data loading and processing pipeline while easily integrating into existing PyTorch applications. FFCV offers a custom file format that can store arbitrary data modalities and enable non-sequential data reading. It stores all samples in a single file but splits the dataset into pages, which trades space for more efficient data reads. If the dataset can fit into memory, it will be cached, and every read of a sample after the first will be read from memory instead of the underlying storage system. In contrast, if the dataset does not fit into memory, FFCV enables efficient data loading by using the PRNG to prefetch data samples. In addition, it uses quasi-random data shuffling in which a permutation of the data samples is loaded into a memory buffer large enough for pages of the dataset. The model then selects from batches existing within the buffer. This reduces the number of accesses to the underlying storage system. FFCV aims to eliminate unnecessary memory allocations by having operations in the data loading and processing pipeline declare memory requirements so that memory allocation occurs only once before each epoch. It also uses just-in-time (JIT) compilation, where possible, during the data processing pipeline to convert Python code to machine code. Since the machine code is no longer under the constraints of the Python interpreter, it enables multi-threading, which by extension enables collaborative reading and writing to and from memory instead of expensive primitives, i.e., threads can share batches of data. Threads also enable data preparation (data copying and augmentations) to be run in parallel as they share the same CUDA context.
@@
Deep Lake (Hambardzumyan et al., 2022) is a data lake for deep learning. It can be used with object storage systems such as AWS S3 (Amazon Web Services, nd), Google Cloud Storage (GCS) (Google Cloud, nd), POSIX file systems, or local in-memory storage. Deep Lake has a Tensor Query Language that extends the SQL Parser from Hyrise (Dreseler et al., 2019). This enables multidimensional array operations and deeper integration with the data lake, including indexing, slicing of arrays, and a large set of convenience functions. There is also a data visualizer made available through a web interface, allowing for the visualization of large-scale data. The Deep Lake streaming data loader delegates data fetching to C++ processes to avoid Python’s GIL. Deep Lake also has a version control system that keeps track of different branches of the dataset in the same storage. This allows queries to be written which specify the dataset version, allowing for more granular and depthful searches. Deep Lake also is capable of storing non-structured data and metadata in a columnar format resulting in fast data streaming speeds.
@@
With the growing size of AI models, efficient checkpointing methods are needed to ensure scalability. Nicolae et al. (et al, 2020e) propose DeepFreeze, an asynchronous checkpointing technique built using Very Low Overhead Checkpoint-Restart (Veloc) (et al, 2019c). Veloc is a lightweight concurrency-optimized checkpointing library aimed at delivering high-performance checkpointing on top of heterogeneous storage hierarchies. DeepFreeze is a transparent checkpointing solution for synchronous data parallel training with layer-wise model parallelism. To increase checkpointing speeds, an asynchronous multilevel checkpointing approach is used in which a local copy of the learning model is captured and persisted in lightweight (i.e., node-local storage of neighboring nodes) and heavy (i.e., PFS) storage to block training for the least amount of time possible. To reduce the overhead of model serialization, a compact binary format is used that leaves out details such as tensor labels. The model’s layers are sliced into shards equal to the number of MPI ranks. The ranks then write checkpoints for their partition of the model to local storage. This allows for models to be checkpointed in parallel, avoiding the stragglers problem. Anthony et al. (Anthony and Dai, 2021) also propose local process model snapshots, which are then stored in node-local storage. Their method uses the SCR-Exa library (Dai, 2022), which is a scalable checkpointing tool for HPC applications. There are two restart methods (cold and hot) available. A cold restart attempts to restart the application using the checkpointing cache within the same allocation of nodes. The hot restart tries to replace faulty nodes with idle spare nodes. Check-N-Run (et al, 2022a) is a checkpointing system for deep-learning recommendation models developed by Facebook. It uses two techniques to address size and bandwidth: differential checkpointing, where only a fraction of the model changes per iteration, and quantization techniques to reduce checkpoint size. The checkpoint system is implemented under the PyTorch framework. Check-N-Run keeps training accuracy as a priority as they state an accuracy loss as small as is unacceptable. Maurya et al. (Maurya et al., 2024) introduce DataStates-LLM, an asynchronous multi-level checkpointing approach to speed up LLM checkpointing. By exploiting the periods of time when tensors and optimizer state shards are immutable, the strategy lazily copies shards of the model from the GPU during periods of immutability into a single host and streams the partial checkpoints asynchronously to persistent storage. The copying from GPU to host and the transfer from host to persistent storage can occur in parallel.
@@ 7. Future Research Directions of I/O for ML
While the ML surge is rushing forward driven by rapidly increasing data collection and application use cases, data management and I/O software have been significantly lagging, especially in the HPC domain. In this section, we identify numerous data management and I/O areas that need to be improved to ensure continued progress and innovation in the training and deployment of ML models.
@@
When preparing data from remote storage locations such as data warehouses, it may be necessary to store intermediate forms of the dataset after preprocessing. Further research could involve an I/O middleware that transparently stores intermediate forms of the dataset in clustered memory or the underlying storage system, depending on the dataset size. By extending the memory to the underlying storage system, additional questions arise, such as how data movement between memory and storage can be optimized for high throughput and low latency and what caching strategies should be used to ensure efficient data access.
@@
With the increasing amount of data for machine learning, ensuring data quality is essential for model accuracy (et al, 2020c). A step towards ensuring high-quality data is to use data warehouses because they often provide data governance tools. However, scientists who use HPC systems for high I/O throughput could take a significant performance hit if limited by the bandwidth of the external network. This motivates the need for tools that can aid in obtaining and monitoring high-quality datasets. These tools could offer data management services, such as integrity constraints, versioning, provenance, and quality assessments, while being scalable and supporting multiple data modalities. A survey on metrics for AI data readiness (Hiniduma et al., 2024) exposes a lack of metrics focusing on quality, bias, and FAIR (Findability, Availability, Interoperability, and Reusability) principle compliance.
@@
Due to the rise of cloud-based storage being used by ML workloads, methods of efficiently streaming data from multiple data sources are needed. In addition, multi-modal models may need to stream data from several data sources, each with varying data modalities. Future research could involve an I/O middleware capable of concurrently optimizing I/O from multiple data sources. The middleware would take into account the varying latencies between the data sources and use methods such as sample caching and sample prefetching to ensure optimal performance. This would involve a cache eviction strategy in which eviction is based on the dynamic latency of the data sources. In addition, prefetch scheduling could also occur dynamically, ensuring that samples from high-latency data sources have a higher priority than low-latency data sources. It would also take into account the sample size when performing these optimizations to ensure the optimal balance between latency and data size.
@@
GPUs have become increasingly popular due to the rise of ML. This is due to their high level of parallelism when doing mathematical computations and, in particular, matrix operations (Steinkraus et al., 2005). This is possible because GPUs, unlike CPUs, contain thousands of cores. This has led to the development of ML domain-specific processing units, such as tensor processing units (TPU)s by Google (et al, 2017). These are specialized processors specifically built for ML workloads. With the wide range of hardware available, future research is needed to determine how data movement and storage techniques should be used effectively across the diverse range of hardware available. Questions for future research include where to place and process samples to prioritize processing units and exploit their contrasting abilities effectively, how to efficiently move data between the diverse range of processing units, and how to provide transparent and reasonable solutions for developers facing these challenges.
@@
As shown in Fig. 1, HPC I/O and ML I/O stacks have been organically evolving for faster implementation of the ML algorithms although there is significant scope for higher performance. As the ML algorithms settle, improving the efficiency of data access in running various applications both on HPC systems as well as cloud computing environments need to take priority. ML application development should also target obtaining portability and efficiency on different platforms. Achieving this will require efforts in benchmarking, evaluation of performance with various programming tools, and developing platform-specific optimization strategies.
Efficient Model Offloading
Model offloading is used in memory-constrained environments to move subsets of the model to lower storage tiers. This has been a prominent area of research in edge computing, where memory constraints are a common limiting factor. With the growth in size of LLMs, many cannot be held in distributed memory. One example of such a model is the DOE’s recently announced one-trillion-parameter (i.e., 4TiB) AI system named “AuroraGPT” (Thakur, 2024). Enabling such a model in a memory-constrained environment requires model offloading. Compounding the issue is that some inference configurations (e.g., ensemble models or pipeline models) require multiple models to be loaded and unloaded from device memory to respond to a single query. Moving the model effectively through the storage tiers to minimize inference response times is a challenging I/O and data management problem where further R&D is needed.
@@ 8. Conclusion
Data plays a pivotal role in machine learning, impacting the learning process and significantly affecting model performance and accuracy. With a wide array of file formats available, each offering distinct I/O optimizations, machine learning applications must carefully select the format that best suits their needs. Additionally, the modality of the dataset influences the choice of file formats, affecting how datasets are accessed and stored throughout the ML lifecycle.
@@
Modern ML frameworks commonly offer several options to optimize data loading and processing, such as parallel data preparation, dataset streaming, and sample prefetching, to optimize the processing and loading of batches of samples. However, caching solutions that work with datasets that do not fit into node local memory are not widely available in ML frameworks such as PyTorch and TensorFlow, often leading to sample accesses always reaching the underlying storage system, which may lead to an I/O bottleneck. Recent advances toward minimizing I/O time include specialized shuffling techniques, I/O schedulers, and domain-specific file formats.
@@ Acknowledgements.
- Amazon Web Services (2024a) Amazon Web Services. 2024a. 1000 Genomes Project. https://registry.opendata.aws/1000-genomes/ Accessed: 2024-09-12.
- Amazon Web Services (2024b) Amazon Web Services. 2024b. AWS Public Datasets. https://registry.opendata.aws/ Accessed: 2024-09-12.
@@
- Apache Software Foundation (2024a) Apache Software Foundation. 2024a. Apache Spark. https://spark.apache.org/.
- Apache Software Foundation (2024b) Apache Software Foundation. 2024b. Hadoop. https://hadoop.apache.org
@@
- Brewer et al. (2020a) Wesley Brewer, Greg Behm, Alan Scheinine, Ben Parsons, Wesley Emeneker, and Robert P. Trevino. 2020a. iBench: a Distributed Inference Simulation and Benchmark Suite. In 2020 IEEE High Performance Extreme Computing Conference (HPEC). 1–6. https://doi.org/10.1109/HPEC43674.2020.9286169
- Brewer et al. (2020b) Wesley Brewer, Greg Behm, Alan Scheinine, Ben Parsons, Wesley Emeneker, and Robert P. Trevino. 2020b. Inference Benchmarking on HPC Systems. In 2020 IEEE High Performance Extreme Computing Conference (HPEC). 1–9. https://doi.org/10.1109/HPEC43674.2020.9286138
@@
- Byna et al. (2020) S. Byna, M. Breitenfeld, B. Dong, Q. Koziol, Elena Pourmal, D. Robinson, Jérome Soumagne, Houjun Tang, Venkatram Vishwanath, and R. Warren. 2020. ExaHDF5: Delivering Efficient Parallel I/O on Exascale Computing Systems. JCST 35 (2020), 145–160.
@@
- Choi et al. (2023) Hyeonseong Choi, Byung Hyun Lee, Se Young Chun, and Jaehwan Lee. 2023. Towards accelerating model parallelism in distributed deep learning systems. PLoS One 18, 11 (Nov. 2023), e0293338.
- Community (2024) Stack Overflow Community. 2024. Search Results for Distributed-Training: Search query for distributed-training up to 2015: distributed training created:..2015 and up to 2024: distributed training created:..2024.
@@
- Dantas et al. (2022a) Marco Dantas et al. 2022a. Accelerating Deep Learning Training Through Transparent Storage Tiering. In 2022 22nd IEEE International Symposium on Cluster, Cloud and Internet Computing (CCGrid). 21–30. https://doi.org/10.1109/CCGrid54584.2022.00011
- Dantas et al. (2022b) Marco Dantas et al. 2022b. Accelerating Deep Learning Training Through Transparent Storage Tiering. In 2022 22nd IEEE International Symposium on Cluster, Cloud and Internet Computing (CCGrid). 21–30. https://doi.org/10.1109/CCGrid54584.2022.00011
@@
- Dau (2018) Hoang Anh et al Dau. 2018. The UCR Time Series Classification Archive. https://www.cs.ucr.edu/~eamonn/time_series_data_2018/.
@@
- et al (2023a) Audibert Andrew et al. 2023a. tf.data service: A Case for Disaggregating ML Input Data Processing. In Proceedings of the 2023 ACM Symposium on Cloud Computing (Santa Cruz, CA, USA). Association for Computing Machinery, New York, NY, USA. https://doi.org/10.1145/3620678.3624666
- et al (2008) Abadi Daniel J. et al. 2008. Column-stores vs. row-stores: how different are they really?. In Proceedings of the 2008 ACM SIGMOD International Conference on Management of Data. Association for Computing Machinery, New York, NY, USA. https://doi.org/10.1145/1376616.1376712